A rod of length 1 =10 cm and the ring radius R=5 cm. Each has a density charge 2= 106 C/m.
Calculate the electric field at a point P and electrostatic force if in P Q=+10-7C.
1
Expert's answer
2021-07-13T10:42:35-0400

"E=E_1+E_2=kQ(\frac 1{l^2}+\frac{1}{R^2})=0.45~\frac{MN}{C},"

"F=F_1+F_2=\frac{kQ\sigma l}{l^2}+\frac{kQ\sigma 2\pi R}{R^2}=kQ\sigma (\frac 1 l+\frac{2\pi}{R})=122~GN."